# Mixture of Scope Experts at Test: Generalizing Deeper Graph Neural Networks with Shallow Variants

### v3

@@ 1 Introduction
Figure 1: Performance on amazon-ratings. (Left) Deeper GNNs exhibit significant performance disparities across different homophily ratios, with the shaded area indicating the homophily distribution. (Right) Deeper GNNs and their shallow variants show a shift in generalization preference across homophily ratios, with the red dotted line indicating the average training set homophily (0.38).
Graph neural networks (GNNs) are emerging as powerful tools for graph mining applications, such as social recommendations [17], traffic prediction [28], and fraud detection [50]. GNNs’ superior performance is largely attributed to graph homophily, where similar nodes tend to be connected. However, some graphs exhibit heterophily, where connected nodes tend to be dissimilar. When aggregating heterophilous information, GNNs tend to generate similar embeddings for nodes of different classes, leading to suboptimal performance.
To find homophily on graphs where heterophily dominates, GNNs need to search for neighbors from higher hops. Typically, a GNN with $L$ layers of propagation can perceive the entire $L$-hop neighborhood for each node, making the scope (i.e., receptive field) size tightly coupled with the GNN’s depth. However, a persistent challenge is that GNNs experience performance degradation as they become deeper. Studies have shown that deeper GNNs suffer from three key issues: reduced expressivity due to over-smoothing [33, 73], model degradation from optimization challenges [81, 42], and large generalization gaps caused by overfitting [74]. While many techniques have been proposed to address the first two issues, these solutions often sacrifice generalization [13], resulting in limited overall performance improvements.
Existing theoretical analyses of deeper GNN generalization adopt a purely global perspective [13, 65], which overlooks the diversity of local structural patterns (e.g., homophily) that appear in real-world graphs. For example, Figure 1 (left) shows that variations in homophily can induce substantial performance discrepancies for deeper GNNs. Recent work has moved toward more realistic assumptions by modeling graphs as mixtures of node subgroups with varying homophily levels [46, 40]. However, these theories are restricted to one- or two-layer GNNs and thus fail to reveal the generalization behavior of GNNs with increasing depths.
@@ 2 Preliminaries
Node Homophily. Homophily metrics are widely used as graph properties to measure the probability of nodes with the same class connected to each other. In this paper, we mainly use a node-wise homophily metric called node homophily [53]. It defines the fraction of neighbors that have the same class for each node $v$: |${\left|\left\{u\in{\mathcal{N}}_{v}:y_{u}=y_{v}\right\}\right|}/{\left|{\mathcal{N}}_{v}\right|}$.
@@ 3.1 Unpacking the Depth Dilemma: Why Do GNNs Struggle with Generalization?
Existing studies have proved that deeper GNNs are more expressive [51, 26, 49, 10]. In practice, however, performance degradation is widely observed when going deep. Unlike previous studies that attribute the failure of deeper GNNs to a single cause, we argue that this failure stems from multiple factors, varying based on GNN architectures. To analyze each factor separately, we break down the error of deeper GNNs on unseen samples. Let $\mathcal{F}$ denote the function class for a given deeper GNN architecture, and $f_{S}\in\mathcal{F}$ denote a function learned on the training set $S$. The population risk (true risk) of $f_{S}$ over the entire data distribution can be decomposed as:
$R\left(f_{S}\right)=$ $\underbrace{\hat{R}_{S}\left(f_{S,ERM}\right)}_{\text{representation error}}+\underbrace{\hat{R}_{S}\left(f_{S}\right)-\hat{R}_{S}\left(f_{S,ERM}\right)}_{\text{optimization error }}+\underbrace{R\left(f_{S}\right)-\hat{R}_{S}\left(f_{S}\right)}_{\text{generalization error }},$ (1)
where $\hat{R}_{S}$ is the empirical risk (training error) and $f_{S,ERM}$ is the empirical risk minimizer. In the following, we discuss these errors and connect them with recent findings: (1) Representation error measures the minimum error over the function class $\mathcal{F}$ on $S$, evaluating the expressive power of the GNN architecture. A major factor limiting deeper GNNs’ expressivity is over-smoothing [33], where node representations become indistinguishable after multiple GNN layers. Recent studies [81] show that GNNs without non-linearity between propagation layers (e.g., SGC) tend to suffer from over-smoothing. (2) Optimization error depicts how well $f_{S}$ trained by calculating the risk difference between the learned model $f_{S}$ with the empirical risk minimizer $f_{S,ERM}$. Research shows that deeper GNNs encounter significant training difficulties [42] and exhibit model degradation [81] as depth increases, leading to decreased accuracy in both training and testing. The primary cause is gradient-related issues such as vanishing gradients [31] and gradient instability [13]. These problems can be mitigated through optimization tricks like skip-connections. (3) Generalization error measures how well the trained model $f_{S}$ generalizes from the training set $S$ to the true distribution. The overfitting [74, 13] issue has been identified to explain why well-trained deeper GNNs with higher expressivity failed to outperform their shallow variants. To address this issue, existing works typically reduce GNN parameters or employ regularization techniques.
@@ 3.2 Subgroup Generalization Bound for GNNs with Varying Scopes: A Data-Centric Perspective
To investigate this question, we conduct a theoretical analysis of how GNNs generalize at different depths. Rather than focusing on model architecture [13], we take a data-centric perspective to understand generalization across diverse local structural patterns. Recent work [46, 40] has shown that real-world graphs comprise node subgroups with varying homophily levels, causing GNNs to exhibit generalization discrepancies across these subgroups. However, these studies only focus on simplified GNNs with one or two layers, which fails to explain deeper GNNs’ failure. Though deeper GNNs have superior expressivity and may generalize better on specific subgroups, these improvements often get obscured by overall performance degradation. To rigorously analyze generalization disparity with respect to GNN depth and examine the role of homophily, we derive a new generalization bound for multi-layer GNNs by extending [46]’s non-i.i.d. PAC-Bayesian analysis on GNNs with one-hop aggregation. Following the assumptions used in [46], we adopt the contextual stochastic block model (CSBM) with 2 classes for a controlled study, which is widely used for graph analysis [4, 68].
@@
The following theorem is based on the PAC-Bayes analysis with margin loss [45, 12]. We aim to bound the generalization gap between the expected loss ${\mathcal{L}}^{0}_{m}(\theta)$ of a test subgroup $m$ for 0 margins and the empirical loss ^$\widehat{{\mathcal{L}}}^{\gamma}_{S}(\theta)$ on train subgroup $S$ for a margin $\gamma$.
@@ 4.1 The MoE Workflow
Scope Experts Training. Different depth GNN models serving as experts of their corresponding scope is a hot topic in recent Graph MoE research [62, 79, 76]. For decoupled GNNs like SGC, we consider the number of propagation layers as depth while keeping the transformation layers fixed. Formally, given a GNN architecture $\mathcal{M}$ and the maximum number of layers $L_{\text{max}}$, we independently train $L_{\text{max}}+1$ models $\{\mathcal{M}^{0},...,\mathcal{M}^{L_{\text{max}}}\}$ with depth from 0 to $L_{\text{max}}$, where $\mathcal{M}^{0}$ is an MLP.
@@
When an expert lacks expressivity, it can suffer from over-smoothing as depth increases. To help identify when over-smoothing happens in experts, we augment their logits with structural encoding, incorporating node smoothness and centrality. Let $\mathbf{X}^{(L)}=\widetilde{{\mathbf{A}}}^{L}{\mathbf{X}}$ represent node features smoothed within the scope size $L$. We measure its distance from both the original feature ${\mathbf{X}}^{(0)}={\mathbf{X}}$ and the final smoothed feature $\mathbf{X}^{(\infty)}=\widetilde{{\mathbf{A}}}^{\infty}{\mathbf{X}}$, where $\widetilde{\mathbf{A}}_{u,v}^{\infty}={\left(d_{u}+1\right)^{\frac{1}{2}}\left(d_{v}+1\right)^{\frac{1}{2}}}/{\left(2|{\mathcal{E}}|+|{\mathcal{V}}|\right)}$ [33]. For each node $v$, we calculate two distance scalars $\bar{\epsilon}^{(L)}_{v}=\|{\mathbf{X}}^{(L)}_{v}-{\mathbf{X}}^{(0)}_{v}\|_{2}$ and $\tilde{\epsilon}^{(L)}_{v}=\|{\mathbf{X}}^{(L)}_{v}-{\mathbf{X}}^{(\infty)}_{v}\|_{2}$ to measure smoothness. We also incorporate PageRank centrality $\pi_{v}$ to encode node position. The structural encoding can be represented as $\bm{\xi}_{\text{struc}}^{(L)}=\left[\bar{\bm{\epsilon}}^{(L)}~{}\|~{}\tilde{\bm{\epsilon}}^{(L)}~{}\|~{}\bm{\pi}\right]$. Finally, we combine these components to form the overall augmented logits for expert $\mathcal{M}^{L}$ as $\bm{\zeta}^{(L)}=\left[\mathbf{Z}^{(L)}~{}\|~{}\bm{\xi}_{\text{label}}^{(L)}~{}\|~{}\bm{\xi}_{\text{struc}}^{(L)}\right]$, where $\bm{\zeta}^{(L)}\in\mathbb{R}^{|{\mathcal{V}}|\times F_{\text{aug}}}$ and $F_{\text{aug}}=C+L_{\text{max}}C+3$.
@@ 4.2 Discussion and Analysis
Runtime analysis. Moscat is a lightweight module with relatively small training time. Please refer to Appendix G.2 for detailed analysis.
@@ 5 Experiments
In this section, we aim to answer the following questions to verify the effectiveness of Moscat. Q1: How does Moscat perform in improving GNNs with varying architectures? Q2: How does Moscat compare to other techniques in enhancing deeper GNNs? Q3: How does each component of Moscat contribute to its overall performance? To understand why Moscat is effective, we further conduct a case study at the end of this section and defer other in-depth experimental analyses to the Appendix: Appendix G.6 investigates how expert training affects Moscat performance, Appendix G.7 provides a visual interpretation of how Moscat enhances GNN generalization, and Appendix G.9 analyzes Moscat’s learned gating weights of different scope experts.
@@ 5.1 Experimental setup
Datasets. We evaluate Moscat on real-world datasets with varying node homophily ratios and graph sizes. We exclude commonly used small heterophilous graphs (fewer than thousands of nodes) due to their high variance and lack of statistical significance [54]. Please see Appendix F.1 for details.
@@
Baselines and Hyperparameters. We compare ours against popular homophilous GNNs, heterophilous GNNs, Graph Transformers, Graph MoEs, and also deeper GNNs with various skip-connections. For baseline hyperparameters, we follow their original papers when available; otherwise, we conduct hyperparameter search using Optuna [2]. For Moscat, we tune only the gating model’s hyperparameters, while experts inherit the same settings as their baseline counterparts. Detailed baseline descriptions and hyperparameters are listed in Appendix F.3 and Appendix F.4.
@@
Squirrel snap-patents arxiv-year Flickr amazon-ratings Penn94 genius ogbn-arxiv Avg.Improv.
@@ 5.2 Performance comparison
Compare with techniques for deeper GNNs. To answer Q2, we compare Moscat with comprehensive skip-connection methods for deeper GCN. We examine (1) residual connections (Res: ResNet-like residual connection [24], II: GCNII-like initial residual, Cat: GraphSAGE-like self concatenation [23]) and (2) learnable gating mechanisms (JK: JKNet with concatenation layer aggregation, Attn: GAMLP-like layer aggregation, G2: gradient gating framework [58]). As shown in Table 3, GCN-Moscat outperforms other skip-connection techniques on Penn94. Additionally, Moscat achieves significant improvements over GNNs with skip-connections on amazon-ratings. However, skip-connections mix output from different scopes, causing experts to make similar predictions and limiting Moscat’s effectiveness. For experiments showing how Moscat performance varies with depth, please refer to Appendix G.3.
@@ A.1 Graph neural networks meet heterophily
GNNs were initially designed under the homophily assumption and have recently gained significant interest due to their superior performance and small parameterization. Various aspects of GNNs have been widely studied, including scalability [23, 77, 5, 20], expressivity [70, 39, 52], and generalization [13, 75, 46].
To extend GNNs to heterophilous graphs, existing works primarily focus on improving higher-order neighborhood utilization [72]. MixHop [1] extracts features from multi-hop neighborhoods in each layer. GCNII [8] prevents over-smoothing in deeper GCNs by proposing initial residual connections and identity mapping. To adapt to graphs with different label patterns, GPR-GNN [11] learns signed scalar weights for the propagated features with different propagation steps. Other works focus on mapping topology [37, 34, 41], using global attentions [15], or exploiting edge directionality [57] to improve learning on heterophilous graphs.

### v4

@@ 1 Introduction
Figure 1: Performance on amazon-ratings. (Left) Deeper GNNs exhibit performance disparities across node subgroups with different homophily ratios, with the shaded area indicating the node distribution. (Right) Deeper GNNs and their shallow variants show a shift in generalization preference across homophily ratios, with the red dotted line indicating the average training set homophily (0.38).
Graph neural networks (GNNs) are emerging as powerful tools for graph mining applications, such as social recommendations [rs-survey], traffic prediction [traffic-survey], and fraud detection [fraud-survey]. GNNs’ superior performance is largely attributed to graph homophily, where similar nodes tend to be connected. However, some graphs exhibit heterophily, where connected nodes tend to be dissimilar. When aggregating heterophilous information, GNNs tend to generate similar embeddings for nodes of different classes, leading to suboptimal performance.
To find homophily on graphs where heterophily dominates, GNNs need to search for neighbors from higher hops. Typically, a GNN with $L$ layers of propagation can perceive the entire $L$-hop neighborhood for each node, making the scope (i.e., receptive field) size tightly coupled with the GNN’s depth. However, a persistent challenge is that GNNs experience performance degradation as they become deeper. Studies have shown that deeper GNNs suffer from three key issues: reduced expressivity due to over-smoothing [over-smoothing, twosides], model degradation from optimization challenges [degradation, training-matters], and large generalization gaps caused by overfitting [revisit-oversmoothing]. While many techniques have been proposed to address the first two issues, these solutions often sacrifice generalization [provable], resulting in limited overall performance improvements.
Existing theoretical analyses of deeper GNN generalization adopt a purely global perspective [provable, wang2024manifold], which overlooks the diversity of local structural patterns (e.g., homophily) that appear in real‑world graphs. For example, Figure 1 (left) shows that variations in homophily can induce substantial performance discrepancies for deeper GNNs. Recent work has moved toward more realistic assumptions by modeling graphs as mixtures of node subgroups with varying homophily levels [demystifying, performance-discrepancy]. However, these theories are restricted to one- or two-layer GNNs and thus fail to reveal the generalization behavior of GNNs with increasing depths.
@@ 2 Preliminaries
Node Homophily. Homophily metrics are widely used as graph properties to measure the probability of nodes with the same class connected to each other. In this paper, we mainly use a node-wise homophily metric called node homophily [geomgcn]. It defines the fraction of neighbors that have the same class for each node $v$: ${\left|\left\{u\in{\mathcal{N}}_{v}:y_{u}=y_{v}\right\}\right|}/{\left|{\mathcal{N}}_{v}\right|}$.
@@ 3.1 Unpacking the Depth Dilemma: Why Do GNNs Struggle with Generalization?
Existing studies have proved that deeper GNNs are more expressive [oono2019graph, huang2020tackling, morris2019weisfeiler, chen2020can]. In practice, however, performance degradation is widely observed when going deep. Unlike previous studies that attribute the failure of deeper GNNs to a single cause, we argue that this failure stems from multiple factors, varying based on GNN architectures. To analyze each factor separately, we break down the error of deeper GNNs on unseen samples. Let $\mathcal{F}$ denote the function class for a given deeper GNN architecture, and $f_{S}\in\mathcal{F}$ denote a function learned on the training set $S$. The population risk (true risk) of $f_{S}$ over the entire data distribution can be decomposed as:
[…]
where $\hat{R}_{S}$ is the empirical risk (training error) and $f_{S,ERM}$ is the empirical risk minimizer. In the following, we discuss these errors and connect them with recent findings: (1) Representation error measures the minimum error over the function class $\mathcal{F}$ on $S$, evaluating the expressive power of the GNN architecture. A major factor limiting deeper GNNs’ expressivity is over-smoothing [over-smoothing], where node representations become indistinguishable after multiple GNN layers. Recent studies [degradation] show that GNNs without non-linearity between propagation layers (e.g., SGC) tend to suffer from over-smoothing. (2) Optimization error depicts how well $f_{S}$ trained by calculating the risk difference between the learned model $f_{S}$ with the empirical risk minimizer $f_{S,ERM}$. Research shows that deeper GNNs encounter significant training difficulties [training-matters] and exhibit model degradation [degradation] as depth increases, leading to decreased accuracy in both training and testing. The primary cause is gradient-related issues such as vanishing gradients [deepgcns] and gradient instability [provable]. These problems can be mitigated through optimization tricks like skip-connections. (3) Generalization error measures how well the trained model $f_{S}$ generalizes from the training set $S$ to the true distribution. The overfitting [revisit-oversmoothing, provable] issue has been identified to explain why well-trained deeper GNNs with higher expressivity failed to outperform their shallow variants. To address this issue, existing works typically reduce GNN parameters or employ regularization techniques.
@@ 3.2 Subgroup Generalization Bound for GNNs with Varying Scopes: A Data-Centric Perspective
To investigate this question, we conduct a theoretical analysis of how GNNs generalize at different depths. Rather than focusing on model architecture [provable], we take a data-centric perspective to understand generalization across diverse local structural patterns. Recent work [demystifying, performance-discrepancy] has shown that real-world graphs comprise node subgroups with varying homophily levels, causing GNNs to exhibit generalization discrepancies across these subgroups. However, these studies only focus on simplified GNNs with one or two layers, which fails to explain deeper GNNs’ failure. Though deeper GNNs have superior expressivity and may generalize better on specific subgroups, these improvements often get obscured by overall performance degradation. To rigorously analyze generalization disparity with respect to GNN depth and examine the role of homophily, we derive a new generalization bound for multi-layer GNNs by extending [demystifying]’s non-i.i.d. PAC-Bayesian analysis on GNNs with one-hop aggregation. Following the assumptions used in [demystifying], we adopt the contextual stochastic block model (CSBM) with 2 classes for a controlled study, which is widely used for graph analysis [csbm:2, csbm:4].
@@
The following theorem is based on the PAC-Bayes analysis with margin loss [subgroup-bound, pac3]. We aim to bound the generalization gap between the expected loss ${\mathcal{L}}^{0}_{m}(\theta)$ of a test subgroup $m$ for 0 margins and the empirical loss $\widehat{{\mathcal{L}}}^{\gamma}_{S}(\theta)$ on train subgroup $S$ for a margin $\gamma$.
@@ 4.1 The MoE Workflow
Scope Experts Training. Different depth GNN models serving as experts of their corresponding scope is a hot topic in recent Graph MoE research [graphmoe, mowst, damoe]. For decoupled GNNs like SGC, we consider the number of propagation layers as depth while keeping the transformation layers fixed. Formally, given a GNN architecture $\mathcal{M}$ and the maximum number of layers $L_{\text{max}}$, we independently train $L_{\text{max}}+1$ models $\{\mathcal{M}^{0},...,\mathcal{M}^{L_{\text{max}}}\}$ with depth from 0 to $L_{\text{max}}$, where $\mathcal{M}^{0}$ is an MLP.
@@
When an expert lacks expressivity, it can suffer from over-smoothing as depth increases. To help identify when over-smoothing happens in experts, we augment their logits with structural encoding, incorporating node smoothness and centrality. Let $\mathbf{X}^{(L)}=\widetilde{{\mathbf{A}}}^{L}{\mathbf{X}}$ represent node features smoothed within the scope size $L$. We measure its distance from both the original feature ${\mathbf{X}}^{(0)}={\mathbf{X}}$ and the final smoothed feature $\mathbf{X}^{(\infty)}=\widetilde{{\mathbf{A}}}^{\infty}{\mathbf{X}}$, where $\widetilde{\mathbf{A}}_{u,v}^{\infty}={\left(d_{u}+1\right)^{\frac{1}{2}}\left(d_{v}+1\right)^{\frac{1}{2}}}/{\left(2|{\mathcal{E}}|+|{\mathcal{V}}|\right)}$ [over-smoothing]. For each node $v$, we calculate two distance scalars $\bar{\epsilon}^{(L)}_{v}=\|{\mathbf{X}}^{(L)}_{v}-{\mathbf{X}}^{(0)}_{v}\|_{2}$ and $\tilde{\epsilon}^{(L)}_{v}=\|{\mathbf{X}}^{(L)}_{v}-{\mathbf{X}}^{(\infty)}_{v}\|_{2}$ to measure smoothness. We also incorporate PageRank centrality $\pi_{v}$ to encode node position. The structural encoding can be represented as $\bm{\xi}_{\text{struc}}^{(L)}=\left[\bar{\bm{\epsilon}}^{(L)}~\|~\tilde{\bm{\epsilon}}^{(L)}~\|~\bm{\pi}\right]$. Finally, we combine these components to form the overall augmented logits for expert $\mathcal{M}^{L}$ as $\bm{\zeta}^{(L)}=\left[\mathbf{Z}^{(L)}~\|~\bm{\xi}_{\text{label}}^{(L)}~\|~\bm{\xi}_{\text{struc}}^{(L)}\right]$, where $\bm{\zeta}^{(L)}\in\mathbb{R}^{|{\mathcal{V}}|\times F_{\text{aug}}}$ and $F_{\text{aug}}=C+L_{\text{max}}C+3$.
@@ 4.2 Discussion and Analysis
Runtime and space analysis. Moscat is a lightweight module with relatively small training time and memory consumption. Please refer to Appendix G.2 and G.3 for detailed analysis.
@@ 5 Experiments
In this section, we aim to answer the following questions to verify the effectiveness of Moscat. Q1: How does Moscat perform in improving GNNs with varying architectures? Q2: How does Moscat compare to other techniques in enhancing deeper GNNs? Q3: How does each component of Moscat contribute to its overall performance? To understand why Moscat is effective, we further conduct a case study at the end of this section and defer other in-depth experimental analyses to the Appendix: Appendix G.7 investigates how expert training affects Moscat performance, Appendix G.8 provides a visual interpretation of how Moscat enhances GNN generalization, and Appendix G.10 analyzes AH’s learned gating weights of different scope experts.
@@ 5.1 Experimental setup
Datasets. We evaluate Moscat on real-world datasets with varying node homophily ratios and graph sizes. We exclude commonly used small heterophilous graphs (fewer than thousands of nodes) due to their high variance and lack of statistical significance [critical]. Please see Appendix F.1 for details.
@@
Baselines and Hyperparameters. We compare ours against popular homophilous GNNs, heterophilous GNNs, Graph Transformers, Graph MoEs, and also deeper GNNs with various skip-connections. For baseline hyperparameters, we follow their original papers when available; otherwise, we conduct hyperparameter search using Optuna [optuna_2019]. For Moscat, we tune only the gating model’s hyperparameters, while experts inherit the same settings as their baseline counterparts. Detailed baseline descriptions and hyperparameters are listed in Appendix F.3 and Appendix F.4.
@@
Squirrel snap-patents arxiv-year Flickr amazon-ratings Penn94 genius ogbn-arxiv Avg.% Improv.
@@ 5.2 Performance comparison
Compare with techniques for deeper GNNs. To answer Q2, we compare Moscat with skip-connection methods for deeper GCN. We examine (1) residual connections (Res: ResNet-like residual connection [resnet], II: GCNII-like initial residual, Cat: GraphSAGE-like self concatenation [sage]) and (2) learnable gating mechanisms (JK: JKNet with concatenation layer aggregation, Attn: GAMLP-like layer aggregation, G2: gradient gating framework [g2gnn]). As shown in Table 3, GCN-Moscat outperforms other skip-connection techniques on Penn94. Additionally, Moscat achieves significant improvements over GNNs with skip-connections on amazon-ratings. However, skip-connections mix output from different scopes, causing experts to make similar predictions and limiting AH’s effectiveness.
Figure 5: Moscat outperforms classic GNNs (e.g., SGC, GCN) and soft-scoping GNNs (e.g., GAMLP, ACMGCN). GAMLP and ACMGCN adaptively learn the scope of SGC and GCN, respectively.
Performance variation with depths. As depth increases, we further investigate how Moscat improves two state-of-the-art deeper GNNs with cross-layer gating, ACMGCN [gamlp] and GAMLP [acmgcn]. Figure 5 shows that classic GNNs like GCN and SGC often degrade rapidly beyond 2 layers. With gating, GAMLP and ACM-GCN sustain gains up to 4-6 layers. In contrast, GNN-Moscat achieves the best results across all depths and continues to improve through 6–10 layers. Notably, when increasing the maximum layers from 1–10, GNN-Moscat with message-passing architectures (GCN, ACMGCN) can derive 4–10% accuracy improvements (Figure 5, top), whereas GNN-Moscat with decoupled architectures (SGC, GAMLP) yields only 1–2% gains (Figure 5, bottom). This gap likely arises from the higher expressivity of message-passing models, which enables depth-varying experts to produce more diverse predictions. See Appendix G.4 for additional experiments.
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
This work is supported by DEVCOM ARL Army Research Office (ARO) under grants W911NF2220159 and W911NF2320186, and by National Science Foundation (NSF) under grant OAC-2209563. Distribution Statement A: Approved for public release. Distribution is unlimited.
@@ A.1 Graph neural networks meet heterophily
GNNs were initially designed under the homophily assumption and have recently gained significant interest due to their superior performance and small parameterization. Various aspects of GNNs have been widely studied, including scalability [sage, saint, pprgo, influence], expressivity [gin, what, lose], and generalization [provable, pmlp, demystifying].
To extend GNNs to heterophilous graphs, existing works primarily focus on improving higher-order neighborhood utilization [tedgcn]. MixHop [mixhop] extracts features from multi-hop neighborhoods in each layer. GCNII [gcnii] prevents over-smoothing in deeper GCNs by proposing initial residual connections and identity mapping. To adapt to graphs with different label patterns, GPR-GNN [gprgnn] learns signed scalar weights for the propagated features with different propagation steps. Other works focus on mapping topology [linkx, global, acmgcn], using global attentions [polynormer], or exploiting edge directionality [dirgnn] to improve learning on heterophilous graphs.